A force of 20N acts on a rocket for 350secs, causing the rockets velocity to increase. By how much does the rocket’s momentum increase?
1
Expert's answer
2020-11-27T10:08:14-0500

We can find the magnitude of the rocket’s momentum increase from the impulse-momentum change equation:


"\Delta p =F\Delta t=20\ N\cdot 350\ s=7000\ N\cdot s."

Answer:

"\Delta p =7000\ N\cdot s."